# Magnetic field and pressure tuning of the heavy fermion antiferromagnet CePdIn

### v1

@@ Abstract
Single-crystal CePdIn samples were synthesized using the Czochralski method. The constituent elements were melted in stoichiometric proportions using a radio-frequency induction furnace, and this boule was used as the starting material for pulling a crystal using a tetra-arc furnace, yielding properties similar to those reported in Ref. [32]. The temperature dependence of the magnetic susceptibility above 2 K and magnetization as a function of field were measured in a Quantum Design Magnetic Property Measurement System (MPMS) with a helium-3 insert (0.4 K and 1.2 K) or a dilution refrigerator insert (0.1 K). The temperature dependence of the specific heat down to 0.4 K was measured in a Quantum Design Physical Property Measurement System (PPMS) with a 3He insert, using a standard relaxation method. The resistivity both at ambient pressure and under pressures in piston-cylinder pressure cells was measured using the standard four probe method. The measurements from 4 K down to 0.3 K were performed in an Oxford Instruments 3He cryostat with a 9 T magnet, and the high temperature parts of the resistivity above 1.9 K were measured in a Quantum Design PPMS. Ac heat capacity under pressure was measured via an ac calorimetric technique [33]. Pressures in the piston-cylinder cell were determined by measuring the superconducting transition temperature of Pb [34]. Daphne 7373 was used as the pressure-transmitting medium.
@@
Figure 3: (a) Temperature dependence of the resistivity $\rho(T)$ and its temperature derivative $\rho(T)/d$$T$ with the current applied perpendicular to the $c$-axis. The arrows mark the two transitions at $T_{\rm{N}}$ and $T_{\rm{M}}$. (b) Temperature dependence of the specific heat as $C(T)/T$ of CePdIn and LaPdIn. The inset displays a plot of $C/T$ versus $T^{2}$, and the solid line represents a linear fit.
@@
Figure 4: (a) Temperature dependence of the magnetic contribution to the specific heat as $C_{\rm{m}}(T)/T$ on a logarithmic temperature scale, with various magnetic fields applied along the $c$-axis. The inset shows the temperature dependence of the magnetic entropy $S_{\rm{m}}/(T)$ at various magnetic fields. (b) Temperature dependence of the resistivity $\rho(T)$ in various applied $c$-axis magnetic fields. The dashed lines correspond to fitting with Fermi liquid behavior. (c) Temperature derivative of the resistivity $\rho(T)/d$$T$ in various applied $c$-axis magnetic fields. The black arrows in (a) and (c) highlight the evolution of the transition temperatures $T_{\rm{N}}$ and $T_{\rm{M}}$ with field, and the red arrow in (a) marks the hump temperature $T_{\rm{max}}$ in the heat capacity.
Figure 4(a) displays the magnetic contribution to the specific heat, $C_{\rm{m}}/T$ with different magnetic fields applied along the $c$-axis, obtained by subtracting the LaPdIn data from that of CePdIn. With increasing field, both $T_{\rm{N}}$ and $T_{\rm{M}}$ shift to lower temperatures, and for fields above 6 T, the sharp transitions in $C_{\rm{m}}/T(T)$ are replaced by a broad hump that moves to higher temperatures with increasing field, indicating the disappearance of antiferromagnetic order. The inset of Fig. 4(a) shows the temperature dependence of the magnetic entropy under various fields. At 10 K, the entropy reaches approximately 70% of $R$ ln 2. Upon entering the magnetically ordered phase, the entropy is reduced to about 35% of $R$ ln 2, indicating a substantial degree of Kondo hybridization. Under applied magnetic fields, the low-temperature entropy decreases monotonically, showing no enhancement of magnetic fluctuations and thus providing no evidence for a field-induced quantum criticality. A similar suppression occurs in the resistivity measurements [Fig. 4(b) and (c)], where $T_{\rm{N}}$ and $T_{\rm{M}}$ also move to lower temperatures with increasing field. Here, $T_{\rm{N}}$ also vanishes entirely above 6 T, suggesting the suppression of magnetic order, and the low temperature resistivity can be described by Fermi liquid behavior (dashed lines).
@@
Figure 6: Temperature dependence of (a) $\rho(T)$ of sample No. 1, (b) $\rho(T)/d$$T$, and (c) ac specific heat as $C/T(T)$ of sample No. 2, of CePdIn under various applied pressures up to 2.3 GPa, measured in a piston-cylinder pressure cell.
@@
To track the pressure evolution of the magnetic ordering, we examine the low-temperature resistivity $\rho(T)$ and its derivative in Figs. 8(a) and (b), respectively. The $T_{\rm{N}}$ exhibits a non-monotonic pressure dependence, whereby initially $T_{\rm{N}}$ decreases with pressure, reaching $\approx 0.9$ K at 1.5 GPa. Upon further compression, $T_{\rm{N}}$ increases and reaches a maximum at 3.6 GPa. Beyond 3.6 GPa, $T_{\rm{N}}$ decreases weakly and abruptly vanishes between 4.3 and 4.9 GPa. Ac calorimetry measurements $C/T(T)$ in the DAC [Fig. 8(c)] show consistent behaviors, in that $T_{\rm{N}}$ reaches a maximum at 3.8 GPa, and suddenly disappears around 4.7–5.0 GPa. The resistivity reaches its maximum at 2.6 GPa, likely due to enhanced spin fluctuations associated with the suppression of magnetic order below this pressure. Note that the aforementioned merging of $T_{\rm{max1}}$ and $T_{\rm{max2}}$ occur at higher pressure than that the magnetic order disappears.
Figure 8: Low temperature (a) $\rho(T)$ of sample No. 3, the corresponding (b) $\rho(T)/d$$T$, and ac specific heat as $C/T$ of sample No.4, of CePdIn measured under various pressures in a DAC. The arrows in (b) mark the antiferromagnetic transition temperature $T_{\rm{N}}$, and arrow in (c) shows the evolution of $T_{\rm{N}}$ with pressure.
@@
Combining resistivity and ac calorimetry measurements, we construct the pressure-temperature phase diagram of CePdIn, which is shown on a logarithmic temperature scale in Fig. 9. Here, $T_{\rm{max2}}$ shows a slight variation with pressure, while $T_{\rm{max1}}$ initially decreases slightly below 1 GPa, then increases steadily, merging with $T_{\rm{max2}}$ above 6.3 GPa. Below 2.6 GPa (AF1 phase), $T_{\rm{N}}$ is suppressed with increasing pressure At 2.6 GPa, $T_{\rm{N}}$ shows an abrupt enhancement, signaling a transition to a distinct magnetic phase labelled AF2. In the AF2 regime, $T_{\rm{N}}$ varies weakly with pressure before disappearing abruptly near 5 GPa.
@@
CePdIn exhibits a more three-dimensional crystal structure compared to its isostructural counterpart CePdAl, which may result in significantly weaker geometric frustration. Magnetic field studies reveal a gradual suppression of the antiferromagnetic phase in CePdIn, contrasting sharply with the complex temperature-field phase diagram of CePdAl that hosts multiple metamagnetic transitions, together with a possible spin-liquid phase [11, 12]. Notably, such field-induced anomalies are also commonly observed in other frustrated ZrNiAl-type systems such as YbAgGe [18], YbPtIn [17], and CePtPb [42], but are absent in CePdIn. Due to the large number of nearly-degenerate ground states in frustrated systems, they often host multiple field-driven phases, both in the aforementioned intermetallics, as well as insulating quantum antiferromagnets [43], and therefore the lack of such phenomena in CePdIn is consistent with weaker frustration. Near the critical field where AFM order vanishes, the specific heat $C/T(T)$ does not show a zero-temperature divergence, but rather a hump, a response that parallels observations in CePdAl [11], YbPtIn [17], and CePtPb [42]. This hump reflects the Zeeman splitting of the CEF ground-state doublet. Above the critical field, the Zeeman splitting exceeds the magnetic coupling, driving a transition from the antiferromagnetic state to the paramagnetic state and resulting in a first-order phase transition. As the magnetic field enhances the CEF ground-state splitting, the characteristic maximum in the heat capacity shifts to higher temperatures with increasing field.
@@
The AF1 and AF2 phases exhibit strikingly different responses to magnetic fields despite their similar values of $T_{\rm{N}}$. While $T_{\rm{N}}$ in AF1 is strongly suppressed by applied fields, it remains significantly more robust in AF2, highlighting that these phases separated near 2.6 GPa correspond to distinct magnetically ordered states. This is further evidenced by transport behavior, where in the AF2 phase, $\rho(T)$ slightly increases below $T_{\rm{N}}$ [Fig. 11 (b)]. One possibility is that this corresponds to the opening of a gap associated with spin-density wave type magnetism, which might suggest that the AF2 phase is of the itinerant-type. In this case, the reduced frustration could lead to the delocalization of the $4f$ electrons within the magnetically ordered phase [3], in contrast to CePdAl where it is proposed that frustration leads to a Kondo destruction type quantum criticality under pressure in the paramagnetic state [15]. On the other hand, for CePdIn at ambient pressure, no magnetic Bragg peaks were observed with neutron diffraction [47], indicating a significantly reduced ordered moment, which suggests that even in the AF1 phase the Kondo effect significantly quenches the Ce-$4f$ moments. Additional sensitive microscopic probes of the magnetic order such as polarized neutron diffraction or muon-spin relaxation could be important for understanding the nature of the magnetic ordering in CePdIn. Chemical substitution can generate positive chemical pressure, functioning in a manner analogous to hydrostatic pressure. Consistent with this, substituting Pd with Ni reduces the unit-cell volume and suppresses the antiferromagnetic order [29]. Notably, however, no counterpart to the AF2 phase identified in our work has been observed, suggesting that chemical pressure and hydrostatic pressure influence the system in different ways.
@@ VI ACKNOWLEDGMENTS
This work was supported by the National Key R&D Program of China (Grant No. 2022YFA1402200 and No. 2023YFA1406303), the National Natural Science Foundation of China (Grants No. 12174332, No. 12034017, No. W2511006, No. 12222410, No. U23A20580, No. 12350710785, and No. 12204159)

### v2

@@ Abstract
Single-crystal CePdIn samples were synthesized using the Czochralski method. The constituent elements were melted in stoichiometric proportions using a radio-frequency induction furnace, and this boule was used as the starting material for pulling a crystal using a tetra-arc furnace, yielding properties similar to those reported in Ref. [32]. The temperature dependence of the magnetic susceptibility above 2 K and magnetization as a function of field were measured in a Quantum Design Magnetic Property Measurement System (MPMS) with a helium-3 insert (0.4 K and 1.2 K). Magnetization as a function of field at 0.1 K was measured by the Faraday method with a homemade capacitive magnetometer [33], mounted on a dilution refrigerator. The temperature dependence of the specific heat down to 0.4 K was measured in a Quantum Design Physical Property Measurement System (PPMS) with a 3He insert, using a standard relaxation method. The resistivity both at ambient pressure and under pressures in piston-cylinder pressure cells was measured using the standard four probe method. The measurements from 4 K down to 0.3 K were performed in an Oxford Instruments 3He cryostat with a 9 T magnet, and the high temperature parts of the resistivity above 1.9 K were measured in a Quantum Design PPMS. Ac heat capacity under pressure was measured via an ac calorimetric technique [34]. Pressures in the piston-cylinder cell were determined by measuring the superconducting transition temperature of Pb [35]. Daphne 7373 was used as the pressure-transmitting medium.
@@
Figure 3: (a) Temperature dependence of the resistivity $\rho(T)$ and its temperature derivative $\rho(T)/d$$T$ with the current applied perpendicular to the $c$-axis. The arrows mark the two transitions at $T_{\rm{N}}$ and $T_{\rm{M}}$. The error bar reflects the finite width of the transition. (b) Temperature dependence of the specific heat as $C(T)/T$ of CePdIn and LaPdIn. The inset displays a plot of $C/T$ versus $T^{2}$, and the solid line represents a linear fit.
@@
Figure 4: (a) Temperature dependence of the magnetic contribution to the specific heat as $C_{\rm{m}}(T)/T$ on a logarithmic temperature scale, with various magnetic fields applied along the $c$-axis. The inset shows the temperature dependence of the magnetic entropy $S_{\rm{m}}(T)$ at various magnetic fields. (b) Temperature dependence of the resistivity $\rho(T)$ in various applied $c$-axis magnetic fields. The dashed lines correspond to fitting with Fermi liquid behavior. (c) Temperature derivative of the resistivity $\rho(T)/d$$T$ in various applied $c$-axis magnetic fields. The black arrows in (a) and (c) highlight the evolution of the transition temperatures $T_{\rm{N}}$ and $T_{\rm{M}}$ with field, and the red arrow in (a) marks the hump temperature $T_{\rm{max}}$ in the heat capacity.
Figure 4(a) displays the magnetic contribution to the specific heat, $C_{\rm{m}}/T$ with different magnetic fields applied along the $c$-axis, obtained by subtracting the LaPdIn data from that of CePdIn. With increasing field, both $T_{\rm{N}}$ and $T_{\rm{M}}$ shift to lower temperatures, and for fields above 6 T, the sharp transitions in $C_{\rm{m}}/T$ are replaced by a broad hump that moves to higher temperatures with increasing field, indicating the disappearance of antiferromagnetic order. The inset of Fig. 4(a) shows the temperature dependence of the magnetic entropy under various fields. At 10 K, the entropy reaches approximately 70% of $R$ ln 2. Upon entering the magnetically ordered phase, the entropy is reduced to about 35% of $R$ ln 2, indicating a substantial degree of Kondo hybridization. Under applied magnetic fields, the low-temperature entropy decreases monotonically, showing no enhancement of magnetic fluctuations and thus providing no evidence for a field-induced quantum criticality. A similar suppression occurs in the resistivity measurements [Fig. 4(b) and (c)], where $T_{\rm{N}}$ and $T_{\rm{M}}$ also move to lower temperatures with increasing field. Here, $T_{\rm{N}}$ also vanishes entirely above 6 T, suggesting the suppression of magnetic order, and the low temperature resistivity can be described by Fermi liquid behavior (dashed lines).
@@
Figure 6: Temperature dependence of (a) $\rho(T)$ of sample No. 1, (b) $\rho(T)/d$$T$, and (c) ac specific heat as $C/T$ of sample No. 2, of CePdIn under various applied pressures up to 2.3 GPa, measured in a piston-cylinder pressure cell. The error bar in (b) at a selected pressure reflects the finite width of the transition.
@@
To track the pressure evolution of the magnetic ordering, we examine the low-temperature resistivity $\rho(T)$ and its derivative in Figs. 8(a) and (b), respectively. The $T_{\rm{N}}$ exhibits a non-monotonic pressure dependence, whereby initially $T_{\rm{N}}$ decreases with pressure, reaching $\approx 0.9$ K at 1.5 GPa. Upon further compression, $T_{\rm{N}}$ increases and reaches a maximum at 3.6 GPa. Beyond 3.6 GPa, $T_{\rm{N}}$ decreases weakly and abruptly vanishes between 4.3 and 4.9 GPa. Ac calorimetry measurements $C(T)/T$ in the DAC [Fig. 8(c)] show consistent behaviors, in that $T_{\rm{N}}$ reaches a maximum at 3.8 GPa, and suddenly disappears around 4.7–5.0 GPa. The resistivity reaches its maximum at 2.6 GPa, likely due to enhanced spin fluctuations associated with the suppression of magnetic order below this pressure. Note that the aforementioned merging of $T_{\rm{max1}}$ and $T_{\rm{max2}}$ occur at higher pressure than that at which the magnetic order disappears.
Figure 8: Low temperature (a) $\rho(T)$ of sample No. 3, the corresponding (b) $\rho(T)/d$$T$, and ac specific heat as $C/T$ of sample No.4, of CePdIn measured under various pressures in a DAC. The arrows in (b) mark the antiferromagnetic transition temperature $T_{\rm{N}}$, and arrow in (c) shows the evolution of $T_{\rm{N}}$ with pressure. The error bars in (b) and (c) at selected pressures reflect the finite width of the transition.
@@
Combining resistivity and ac calorimetry measurements, we construct the pressure-temperature phase diagram of CePdIn, which is shown on a logarithmic temperature scale in Fig. 9. Here, $T_{\rm{max2}}$ shows a slight variation with pressure, while $T_{\rm{max1}}$ initially decreases slightly below 1 GPa, then increases steadily, merging with $T_{\rm{max2}}$ above 6.3 GPa. Below 2.6 GPa (AF1 phase), $T_{\rm{N}}$ is suppressed with increasing pressure At 2.6 GPa, $T_{\rm{N}}$ shows an abrupt enhancement, signaling a transition to a distinct magnetic phase labelled AF2. In the AF2 regime, $T_{\rm{N}}$ varies weakly with pressure before disappearing abruptly near 5 GPa. Note that the pressure in the DAC was calibrated at room temperature. Upon cooling, a slight increase in pressure is generally expected in a DAC. Since in situ pressure determination at low temperature is not available in the present setup, this effect is not explicitly included in the pressure–temperature phase diagram. Nevertheless, a direct comparison between measurements performed in a piston–cylinder cell and those carried out in a DAC reveals no visible deviation in the evolution of $T_{\rm{N}}$ with pressure (Fig. 9), indicating that the pressure change upon cooling in the DAC is small and does not qualitatively affect the phase diagram.
@@
CePdIn exhibits a more three-dimensional crystal structure compared to its isostructural counterpart CePdAl, which may result in significantly weaker geometric frustration. Magnetic field studies reveal a gradual suppression of the antiferromagnetic phase in CePdIn, contrasting sharply with the complex temperature-field phase diagram of CePdAl that hosts multiple metamagnetic transitions, together with a possible spin-liquid phase [11, 12]. Notably, such field-induced anomalies are also commonly observed in other frustrated ZrNiAl-type systems such as YbAgGe [18], YbPtIn [17], and CePtPb [43], but are absent in CePdIn. Due to the large number of nearly-degenerate ground states in frustrated systems, they often host multiple field-driven phases, both in the aforementioned intermetallics, as well as insulating quantum antiferromagnets [44], and therefore the lack of such phenomena in CePdIn is consistent with weaker frustration. Near the critical field where AFM order vanishes, the specific heat $C/T$ does not show a zero-temperature divergence, but rather a hump, a response that parallels observations in CePdAl [11], YbPtIn [17], and CePtPb [43]. This hump reflects the Zeeman splitting of the CEF ground-state doublet. Above the critical field, the Zeeman splitting exceeds the magnetic coupling, driving a transition from the antiferromagnetic state to the paramagnetic state and resulting in a first-order phase transition. As the magnetic field enhances the CEF ground-state splitting, the characteristic maximum in the heat capacity shifts to higher temperatures with increasing field.
@@
The AF1 and AF2 phases exhibit strikingly different responses to magnetic fields despite their similar values of $T_{\rm{N}}$. While $T_{\rm{N}}$ in AF1 is strongly suppressed by applied fields, it remains significantly more robust in AF2, highlighting that these phases separated near 2.6 GPa correspond to distinct magnetically ordered states. This is further evidenced by transport behavior, where in the AF2 phase, $\rho(T)$ is slightly enhanced below $T_{\rm{N}}$ [Fig. 11 (b)]. One possibility is that this corresponds to the opening of a gap associated with spin-density wave type magnetism, which might suggest that the AF2 phase is of the itinerant-type. In this case, the reduced frustration could lead to the delocalization of the $4f$ electrons within the magnetically ordered phase [3], in contrast to CePdAl where it is proposed that frustration leads to a Kondo destruction type quantum criticality under pressure in the paramagnetic state [15]. On the other hand, for CePdIn at ambient pressure, no magnetic Bragg peaks were observed with neutron diffraction [48], indicating a significantly reduced ordered moment, which suggests that even in the AF1 phase the Kondo effect significantly quenches the Ce-$4f$ moments. Alternatively, strong critical antiferromagnetic fluctuations or a suppression of Kondo coherence could also contribute to an increase in resistivity around $T_{\rm{N}}$ in the AF2. Additional sensitive microscopic probes of the magnetic order such as polarized neutron diffraction or muon-spin relaxation could be important for understanding the nature of the magnetic ordering in CePdIn. Chemical substitution can generate positive chemical pressure, functioning in a manner analogous to hydrostatic pressure. Consistent with this, substituting Pd with Ni reduces the unit-cell volume and suppresses the antiferromagnetic order [29]. Notably, however, no counterpart to the AF2 phase identified in our work has been observed, suggesting that chemical pressure and hydrostatic pressure influence the system in different ways.
@@ VI ACKNOWLEDGMENTS
This work was supported by the National Key R&D Program of China (Grant No. 2022YFA1402200 and No. 2023YFA1406303), the National Natural Science Foundation of China (Grants No. 12174332, No. 12034017, No. W2511006, No. U23A20580, No. 12350710785, and No. 12204159), and the Zhejiang Provincial Natural Science Foundation of China (Grant No. LRG26A040001).